# Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts

### v1

@@ Abstract
The rapid advancement of Vision-Language Models (VLMs) has expanded multimodal applications, yet evaluations often focus on basic tasks like object recognition, overlooking abstract aspects such as personalities and values. To address this gap, we introduce Value-Spectrum, a visual question-answering benchmark aimed at assessing VLMs based on Schwartz’s value dimensions that capture core values guiding people’s beliefs and actions across cultures. We constructed a vectorized database of over 50,000 short videos sourced from TikTok, YouTube Shorts, and Instagram Reels, covering multiple months and a wide array of topics such as family, health, hobbies, society, and technology. We also developed a VLM agent pipeline to automate video browsing and analysis. Benchmarking representative VLMs on Value-Spectrum reveals significant differences in their responses to value-oriented content, with most models exhibiting a preference for hedonistic topics. Beyond identifying natural preferences, we also explored the ability of VLM agents to adopt specific personas when explicitly prompted, revealing insights into the models’ adaptability in role-playing scenarios. These findings highlight the potential of Value-Spectrum as a comprehensive evaluation set for tracking VLM advancements in value-based tasks and for developing more sophisticated role-playing AI agents.
Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts
Jingxuan Li1,*, Yuning Yang1,*, Shengqi Yang2, Yizhou Zhao1, Ying Nian Wu1, 1University of California, Los Angeles, 2Los Alamos National Laboratory, Correspondence: madili@g.ucla.edu
11footnotetext: The authors contributed equally to this work and are listed in alphabetical order by first name.
@@ 1 Introduction
Figure 1: Exploring Value-Driven Role-Playing in Vision-Language Models. This study investigates how VLMs adopt assigned personas to align value traits and preferences within the social media contexts.
@@
Figure 2: Overview of Video Distribution of Value-Spectrum Dataset. We collected an abundance of short videos relevant to 10 Schwartz values. The area of circles centered at each keyword represents the number of relevant videos in the database.
Vision-Language Models (VLMs), which align Large Language Models (LLMs) with pre-trained vision encoders through cross-modal alignment training, have shown impressive perceptual and cognitive capabilities in tasks like visual question answering and image captioning, especially when processing everyday images.  Zhou et al. (2019); Radford et al. (2021); Zhang et al. (2023); Lu et al. (2024) As these abilities advance, VLMs are being deployed globally, bringing to light concerns within the responsible AI community about the potential misalignment of synthetic preferences Li et al. (2024), personalities Serapio-García et al. (2023), and values Ren et al. (2024).
@@
Building on these capabilities, questions arise about whether VLMs, like LLMs, exhibit inherent personality traits or synthetic preferences that align with human-designed values. Addressing this gap, our study explores two key questions: (1) Do VLMs exhibit distinct preference traits? (2) Can we embed a human-designed persona into VLMs to influence their behavior and preferences? To systematically evaluate these questions, we propose a framework that evaluates VLM personality traits through an analysis of their value systems. In human psychology, values are guiding principles that influence attitudes, beliefs, and traits, forming a core component of personality Schwartz (2012). By evaluating how VLMs prioritize these values, we can gain insights into their preference traits and alignment with human-designed personas.
We introduce Value-Spectrum, a benchmark designed to reveal value-related traits in VLMs through value-based keyword searches. Using VLM agents embedded within social media platforms, we compiled a dataset of $50,191$ unique video records from different social media platforms, each segment aligned with distinct value dimensions from the theory of basic human values (see Fig. 2). By presenting the screenshots of these videos to various VLM agents and asking value-based preference questions, we quantified preference differences between models.
@@
Figure 3: VLM Agent for Social Media Video Collection and Interaction. Our pipeline evaluates and stores VLM’s value preferences across social media platforms. This pipeline uses a VLM agent to collect and process videos from social media platforms like Instagram, TikTok, and YouTube. AI models such as GPT, Gemini, Qwen, and Claude interpret the content of the videos. The system then presents a video to the user and asks if the content is interesting. Based on the user’s yes/no response, the agent either continues watching the video or moves to the next one. This creates an interactive system that tailors the video collection experience to user preferences.
@@
Building upon this benchmark, we systematically evaluated several state-of-the-art VLMs to analyze their alignment with value dimensions. Our findings revealed a shared tendency across models: most exhibited a strong inclination towards certain values, such as Hedonism, with consistently positive responses up to 75% of the time. However, alignment varied significantly across models. For instance, CogVLM and Qwen-VL-Plus demonstrated broad and consistent alignment across all value dimensions, while closed-source models like GPT-4o and Claude-3.5 showed distinct preferences, favoring values such as Self-direction and Hedonism. In contrast, Blip-2 exhibited minimal alignment across all value dimensions, highlighting its limitations in expressing stable preferences. These findings highlight diverse capabilities among VLMs in aligning with human value dimensions but also reveal challenges in inducing consistent, human-designed personas across models.
@@
To address these challenges, we propose two strategies—Simple and ISQ—to systematically adjust value decompositions and induce specific personas within VLMs. By evaluating the effectiveness of these strategies across multiple platforms, our experiments show that TikTok serves as an optimal testing environment for inducing VLM personalities, with models demonstrating the strongest alignment under ISQ. Notably, Claude-3.5 achieved the largest gains with ISQ, whereas Blip-2 showed no improvement under either strategy, underscoring fundamental differences in model adaptability. These results validate our approach and provide a pathway for optimizing VLMs’ personality alignment in real-world applications.
@@
  • We present a dataset of over $50$k short videos spanning diverse topics, social media platforms, and release dates, designed to systematically evaluate the personalities and preferences of VLMs.
  • We propose Value-Spectrum, a benchmark for quantifying VLM value preferences, using social media-based assessments to reveal stable traits across different VLM architectures.
  • We embed specific role-play preferences into VLMs using two strategies(simple and ISQ) to adjust value decompositions, achieving improved personality alignment in cross-platform and real-world tests.
@@ 2.1 Vision-Langauge Agents
A Vision-Language Model (VLM) is a fusion of vision and natural language models. Its inputs are images and textual descriptions and it learns to discover the knowledge from the two modalities. The recent development of large VLMs is rapidly advancing the field of AI. These models have the potential to revolutionize various industries and tasks, showcasing their power in plot and table identifying (Liu et al., 2022), visual-question answer (VQA) (Hu et al., 2024), image captioning (Bianco et al., 2023), and e.t.c. Following Niu et al. (2024), the environment for VLM agents to interact with social media can be constructed, we design an automated control pipeline that guides the agent to continuously interact with social networks.
@@ 2.3 Sentiment, Personality, and Value
The community has been using machine learning-based models to study human sentiment (Malviya et al., 2020), personality (Stachl et al., 2020), and value (Qiu et al., 2022). revious studies focused on human personality classification (e.g. Myers-Briggs Type Indicator (MBTI)) and machine behaviors (i.e., LLMs’ personality Serapio-García et al. (2023)). Inspired by recent studies on indirectly revealing AI agent’s personalities by physcogical exams Jiang et al. (2024), Questionnaires Huang et al. (2023), and cultural perspectives Kovač et al. (2023), we use the new perspective of revealing VLM’s persona by examining machine behaviors and personalities to evaluate their performance on mainstream social media platforms.
Figure 4: Schwartz Value-Based Video Retrieval Pipeline. This pipeline retrieves videos based on Schwartz values by associating each value with relevant words, such as linking Hedonism to topics like vacation. These word associations are converted into vector queries, retrieving matching video content from a database.
@@ 3 Data Collection
Inspired by ScreenAgent Niu et al. (2024), our work leverages a VLM-driven graphical user interface (GUI) agent to autonomously navigate popular social media platforms. This agent conducts random walks through social media platforms where it observes and captures video links alongside screenshots. The data collected are stored in a vector database Han et al. (2023), creating a structured repository optimized for value decomposition and efficient retrieval. We aim to analyze VLM behavior across diverse social contexts and reveal VLMs’ preferences. The automated data collection (see Fig 3) process efficiently fetched a large volume of diverse content, enabling the scope and depth of the analysis that traditional manual collection methods could not achieve.
The resulting dataset comprises $50,191$ video links sourced from Instagram ($32\%$), YouTube ($29\%$), and TikTok ($39\%$). Each entry includes the video link, a representative screenshot, and meta-information such as platform name and post date, capturing a comprehensive snapshot of content posted between July 31st, 2024, and October 31st, 2024. By distributing data evenly across these platforms, we enable balanced analysis and facilitate unbiased value decomposition across social media content. This innovative data set empowers researchers to explore the behavior of VLM in a systematic and organized way, fostering deeper insights into model interpretation and the dynamics of social media.
To examine the distribution of video themes in this data set, we take the screenshot of each video in the beginning as the representation of the video content. This selection is justified in Section 6. We then vectorize the image and define its relevance to a specific keyword as the L2 distance. In Figure 2 we present the abundances of videos that are relevant to ten Schwartz Values. Specifically, for each Schwartz Value we populate representative keywords generated from GPT-4o. The area of the transparent circle is proportional to the number of videos that lie within a distance of $1.5$ to the corresponding normalized keyword vector. Through this simple diagram we find that videos relevant to these Schwartz Value Dimensions Achievement, Hedonism, and Power appear most frequently, while videos about Tradition are relatively rare.
@@ 4 Evaluating VLM’s Preferences
Extending the idea of analysing Large Langauge Models’ human likeness (Shanahan et al., 2023; Kovač et al., 2023) to Vision-Language Models (VLMs) with both pictorial and textual inputs, we asks: Do VLMs also exhibit inherent preferences?
To answer this question, We explore a diverse set of VLMs including GPT-4o OpenAI (2023) , Gemini-1.5 Team et al. (2023), Claude-3.5 Anthropic (2023), Qwen-VL-Plus Bai et al. (2023), CogVLM Wang et al. (2023), and Blip-2 Li et al. (2023) to assess their value preferences. We quantified VLMs’ preferences by evaluating their attitudes toward the $10$ Schwartz values: Hedonism, Self-direction, Security, Achievement, Stimulation, Tradition, Benevolence, Universalism, Conformity, and Power. This approach enables us to construct a comprehensive profile of each model’s value preferences and to identify its unique value traits.
@@ 4.1 Preference Retrieval
To evaluate a VLM’s preference for a specific Schwartz value, we collected each model’s responses to images associated with several keywords related to the value (see Figure 4). For instance, we selected the keywords Equality, Globe, and Handshake for the Universalism dimension because they closely align with its core principles of fairness and global awareness. For each keyword linked to the value, each model reviews five images and answers their attitude towards each image.
@@
The answer to the first question is processed into either yes ($1$) or no (0), and the average score is calculated to evaluate the intensity of the model’s preference for a given value (e.g., Universalism), ranging from 00 to $1$.
@@ 4.2 Preference Patterns
(1) Global Pattern: After summarizing the preference score across all VLMs, we found most of them tend to prioritize certain values, such as Hedonism, over others. The results indicate a general preference for Hedonism and Self-direction, while showing less excitement for Conformity and Power. This trend reflects current societal tendencies toward the pursuit of pleasure and personal goal in life Wrosch and Scheier (2003). The specific ranking is presented in Figure 6.
(2) Range Consistency: As shown in Figure 5, each model’s preference scores remain within a narrow band of approximately ±0.1 around a central value. Models display varying levels of engagement with the content: some, like Blip-2, are more reserved, occupying a smaller area on the plot with lower preference scores, while others, like CogVLM, show greater enthusiasm, demonstrating interest across all inputs with higher preference scores.
(3) Individual Model Variations: When analyzed individually, some models, such as Qwen VL Plus, exhibit consistently high preferences across all 10 Schwartz values. In contrast, other models, like Claude, displays more specific and nuanced preferences as indicated in the standard deviation of responses given by the model across all values (Figure 7).
Value GPT4o Q-wen vl plus Blip 2 CogVLM Gemini 1.5 Claude 3.5
@@
Table 1: Comparison of model preferences based on Schwartz’s 10 value. The average value and standard deviation show the values for each model (GPT4o, Q-wen vl plus, Blip 2, CogVLM, Gemini 1.5, and Claude 3.5) across various dimensions are presented to 2 significant figures.
@@
The preference scores across the VLMs highlight varying levels of engagement and unique sentiment patterns. GPT-4o with scores ranging from $0.7$ to $1$, shows strong preferences, particularly for Self-direction and Hedonism. Gemini-1.5 (pro) shows a more neutral perspective with scores hovering around $0.5$, giving its highest score to Achievement and lowest to Power. Claude-3.5 exhibits the most fluctuation, with a notable aversion to Power and a strong preference for Hedonism. Qwen-VL-Plus maintains a high score of around $0.9$ across all dimensions, with a slight aversion to Power but otherwise affirming most content, aligning with the broader trend against Power. CogVLM consistently scores above $0.9$, with a particularly strong preference for Benevolence, setting it apart from others. In contrast, Blip-2 ranks the lowest with scores around $0.2$, showing minimal engagement and offering subjective or passive responses, reflecting its lack of distinct sentiments and preferences.

### v2

# Jingxuan Li UCLA

@@ Abstract
The recent progress in Vision-Language Models (VLMs) has broadened the scope of multimodal applications. However, evaluations often remain limited to functional tasks, neglecting abstract dimensions such as personality traits and human values. To address this gap, we introduce Value-Spectrum, a novel Visual Question Answering (VQA) benchmark aimed at assessing VLMs based on Schwartz’s value dimensions that capture core values guiding people’s preferences and actions. We designed a VLM agent pipeline to simulate video browsing and constructed a vector database comprising over 50,000 short videos from TikTok, YouTube Shorts, and Instagram Reels. These videos span multiple months and cover diverse topics, including family, health, hobbies, society, technology, etc. Benchmarking on Value-Spectrum highlights notable variations in how VLMs handle value-oriented content. Beyond identifying VLMs’ intrinsic preferences, we also explored the ability of VLM agents to adopt specific personas when explicitly prompted, revealing insights into the adaptability of the model in role-playing scenarios. These findings highlight the potential of Value-Spectrum as a comprehensive evaluation set for tracking VLM alignments in value-based tasks and abilities to simulate diverse personas.
[Uncaptioned image] Value-Spectrum: Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts
Jingxuan Li1,*, Yuning Yang1,*, Shengqi Yang2, Linfan Zhang1, Ying Nian Wu1, 1University of California, Los Angeles, 2Los Alamos National Laboratory, Correspondence: madili@g.ucla.edu
@@ 1 Introduction
Figure 1: Exploring Value-Driven Role-Playing in Vision-Language Models. This study investigates how VLMs adopt assigned personas to align value traits and preferences within social media contexts.
@@
Figure 2: Overview of short video screenshots distribution of Value-Spectrum Dataset. We collected an abundance of short video screenshots relevant to 10 Schwartz values. The area of circles centered at each keyword represents the number of relevant videos in the database.
Vision-Language Models (VLMs), built upon Large Language Models (LLMs) with pre-trained vision encoders through cross-modal alignment training, have shown impressive perceptual and cognitive capabilities in tasks like VQA and image captioning Zhou et al. (2019); Radford et al. (2021); Zhang et al. (2023); Lu et al. (2024). Recent research has identified that LLMs exhibit distinct preferences Li et al. (2024), personalities Serapio-García et al. (2023), and values Ren et al. (2024). In addition, some studies have explored the potential of LLMs as role-playing agents to simulate various personas Wang et al. (2023b); Chen et al. (2024a). Questions thus arise about whether VLMs, as visual extensions of LLMs, also exhibit inherent preferences and whether they can be induced to role-play specific personas.
@@
To address these concerns, our study explores two key questions: (1) Do VLMs exhibit preference traits? (2) Could VLMs adapt their traits to role-play specific human-designed personas, aligning their behaviors and preferences to match predefined roles? To answer the questions, we propose a framework that systematically evaluates VLM preference traits through an analysis of their values, i.e., the guiding principles that influence (human) attitudes, beliefs, and traits Schwartz (2012). By evaluating how VLMs prioritize these values, we can gain insights into their preference traits and alignment with human-designed personas.
In this paper, we introduce Value-Spectrum  [^1] The dataset can be downloaded at https://anonymous.value-spectrum.com/, a benchmark designed to systematically evaluate preference traits in VLMs through visual content from social media. Our framework utilizes VLM agents embedded within social media platforms to collect a dataset of $50,191$ unique short video screenshots spanning a wide range of topics, including lifestyle, technology, health, and more. To enable scalable evaluation, we construct a vector database using the CLIP model Radford et al. (2021), facilitating keyword-driven retrieval of images aligned with specific value dimensions. Specifically, we manually curate ten representative keywords for each Schwartz value and retrieve relevant images from the database. See Fig. 2 for examples of keywords and images. These images are then presented to VLMs alongside preference questions designed to probe their alignment with each value dimension.
@@
Figure 3: VLM Agent for Social Media Video Collection and Interaction. Our pipeline uses a VLM agent to collect and process videos from social media platforms like Instagram, TikTok, and YouTube. AI models such as GPT, Gemini, Qwen, and Claude interpret the content of the videos. The system then presents a video to the user and asks if the content is interesting. Based on the user’s yes/no response, the agent either continues watching the video or moves to the next one. This creates an interactive system that tailors the video collection experience to user preferences.
@@
Our findings reveal a shared tendency among models: most exhibit a strong inclination towards certain values, such as Hedonism. However, preferences still varied across models. For instance, CogVLM and Qwen-VL-Plus demonstrate broad and consistent preferences across all value dimensions, while closed-source models like GPT-4o and Claude 3.5 Sonnet show distinct preferences, favoring values like Self-direction over others. In contrast, Blip-2 exhibits minimal preferences across all value dimensions, highlighting its limitations in expressing stable preferences.
@@
In addition to the static preferences of VLMs, we evaluate the ability of VLMs to adapt their inherent preferences to role-play specific personas. We propose two strategies, Simple and ISQ, to assess the effectiveness of different prompt techniques in inducing VLMs with injected persona. By evaluating the effectiveness of these strategies across multiple platforms, our experiments show that TikTok serves as an optimal testing environment for inducing VLM personalities, with models demonstrating the strongest alignment under ISQ. Notably, Claude 3.5 Sonnet achieved the largest gains with ISQ, whereas Blip-2 showed no improvement under either strategy, underscoring fundamental differences in model adaptability.
@@
  • We present a dataset of over $50$k short video screenshots spanning diverse topics, social media platforms, and release dates, designed to systematically evaluate the personalities and preferences of VLMs.
  • We propose Value-Spectrum, a benchmark for quantifying VLM value preferences, using social media-based assessments to reveal stable traits across different VLMs.
  • We embed specific role-play personas into VLMs using two strategies(simple and ISQ) to adjust value traits, achieving improved personality alignment in real-world interactions.
@@ 2.1 Vision-Langauge Agents
Vision-Language Models take inputs as images and textual descriptions, and they learn to discover the knowledge from the two modalities. The recent development of large VLMs is rapidly advancing the field of AI. These models have the potential to revolutionize various industries and tasks, showcasing their power in plot and table identifying (Liu et al., 2022), visual-question answer (VQA) (Hu et al., 2024), image captioning (Bianco et al., 2023), and e.t.c. Following Niu et al. (2024), the environment for VLM agents to interact with social media can be constructed, we design an automated control pipeline that guides the agent to continuously interact with social networks.
@@ 2.3 Sentiment, Personality, and Value
The community has been using machine learning-based models to study human sentiment (Malviya et al., 2020), personality (Stachl et al., 2020), and value (Qiu et al., 2022). previous studies focused on human personality classification (e.g., Myers-Briggs Type Indicator (MBTI)) and machine behaviors (i.e., LLMs’ personality Serapio-García et al. (2023)). Inspired by recent studies on indirectly revealing AI agent’s personalities by physiological exams Jiang et al. (2024), Questionnaires Huang et al. (2023), and cultural perspectives Kovač et al. (2023), we use the new perspective of revealing VLM’s persona by examining machine behaviors and personalities to evaluate their performance on mainstream social media platforms.
Figure 4: Schwartz Value-Based Video Retrieval Pipeline. This pipeline retrieves video screenshots based on Schwartz values by associating each value with relevant words, such as linking Hedonism to topics like vacation. These word associations are converted into vector queries, retrieving matching video content from a database.
@@ 3 Data Collection
Inspired by ScreenAgent Niu et al. (2024), our work leverages a VLM-driven graphical user interface (GUI) agent to autonomously navigate popular social media platforms. This agent conducts random walks through social media platforms where it observes and captures video links alongside screenshots. The data collected are stored in a vector database Han et al. (2023), creating a structured repository optimized for value decomposition and efficient retrieval. We aim to analyze VLM behavior across diverse social contexts and reveal VLMs’ preferences. The automated data collection (see Figure 3) process efficiently fetched a large volume of diverse content, enabling the scope and depth of the analysis that traditional manual collection methods could not achieve.
The resulting dataset comprises $50,191$ video links sourced from Instagram ($32\%$), YouTube ($29\%$), and TikTok ($39\%$). Each entry includes the video link, a screenshot, and meta-information such as platform name and post date, capturing a comprehensive snapshot of content posted between July 31st, 2024, and October 31st, 2024. By distributing data evenly across these platforms, we enable balanced analysis and facilitate unbiased value decomposition across social media content. This innovative data set empowers researchers to explore the behavior of VLM in a systematic and organized way, fostering deeper insights into model interpretation and the dynamics of social media.
To examine the distribution of video themes in this data set, we take a screenshot of each video in the beginning as the representation of the video content. We then vectorize the image and define its relevance to a specific keyword as the L2 distance. In Fig. 2 we present the abundances of videos that are relevant to ten Schwartz Values. Specifically, for each Schwartz value, we populate representative keywords generated from GPT-4o. The area of the transparent circle is proportional to the number of videos that lie within a distance of $1.5$ to the corresponding normalized keyword vector. Through this simple diagram, we find that videos relevant to these Schwartz Value Dimensions Achievement, Hedonism, and Power appear most frequently, while videos about Tradition are relatively rare.
@@ 4 Evaluating VLM’s Preferences
Extending the idea of analyzing LLM’ human likeness (Shanahan et al., 2023; Kovač et al., 2023) to VLMs with both pictorial and textual inputs, we ask: Do VLMs also exhibit inherent preferences?
To answer this question, We explore a diverse set of VLMs including GPT-4o OpenAI (2023), Gemini 1.5 Pro Team et al. (2023), Claude 3.5 Sonnet Anthropic (2023), Qwen-VL-Plus Bai et al. (2023), CogVLM Wang et al. (2023a), and Blip-2 Li et al. (2023) to assess their value preferences. We quantified VLMs’ preferences by evaluating their attitudes toward the $10$ Schwartz values: Hedonism, Self-direction, Security, Achievement, Stimulation, Tradition, Benevolence, Universalism, Conformity, and Power. This approach enables us to construct a comprehensive profile of each model’s value preferences and to identify its unique value traits.
@@ 4.1 Preference Retrieval
To evaluate a VLM’s preference for a specific Schwartz value, we collected each model’s responses to images associated with several keywords related to the value (see Fig. 4). For instance, we selected the keywords Equality, Globe, and Handshake for the Universalism dimension because they closely align with its core principles of fairness and global awareness. For each keyword linked to the value, each model reviews five images and answers their attitude towards each image.
@@
The answer to the first question is processed into either yes ($1$) or no (0), and the average score is calculated to evaluate the intensity of the model’s preference for a given value (e.g., Universalism).
@@ 4.2 Preference Patterns
(1) Global Pattern: After summarizing the preference score across all VLMs, we found most of them tend to prioritize certain values over others. The results indicate a general preference for Hedonism and Self-direction while showing relatively less excitement for Conformity and Power. This trend reflects current societal tendencies toward the pursuit of pleasure and personal goals in life Wrosch and Scheier (2003). The specific ranking is presented in Fig. 6.
(2) Range Consistency: As shown in Fig. 5, each model’s preference scores remain within a narrow band of approximately ±$0.1$ around a central value. Models display varying levels of engagement with the content: some, like Blip-2, are more reserved, occupying a smaller area on the plot with lower preference scores, while others, like CogVLM, show greater enthusiasm, demonstrating interest across all inputs with higher preference scores.
(3) Individual Model Variations: When analyzed individually, some models, such as Qwen-VL- Plus, exhibit consistently high preferences across all 10 Schwartz values. In contrast, other models, like Claude 3.5 Sonnet, display more specific and nuanced preferences as indicated in the standard deviation of responses given by the model across all values (Fig. 7).
Value GPT-4o Qwen-VL-Plus Blip-2 CogVLM Gemini 1.5 Pro Claude 3.5 Sonnet
@@
Table 1: Model preference scores and standard deviation based on Schwartz’s 10 values. Higher scores indicate stronger preferences, while higher standard deviations reflect greater uncertainty. Average scores and standard deviations for each model are reported as two significant figures.
@@
The preference scores across VLMs highlight varying levels of value traits. GPT-4o with scores ranging from $0.7$ to $1$, shows strong preferences for Self-direction and Hedonism. Gemini 1.5 Pro shows a more neutral perspective with scores hovering around $0.5$, giving its highest score to Achievement and lowest to Power. Claude 3.5 Sonnet exhibits the most fluctuation, with a notable aversion to Power and a strong preference for Hedonism. Qwen-VL-Plus maintains a high score of around $0.9$ across all dimensions, with a slight aversion to Power. CogVLM consistently scores above $0.9$, with a particularly strong preference for Benevolence, setting it apart from others. In contrast, Blip-2 ranks the lowest with scores around $0.2$, showing minimal engagement and offering subjective or passive responses, reflecting its lack of distinct sentiments and preferences.